OpenStudy (anonymous):

Write the sum or difference in simplest form. On Monday, Cathey read 1/6 of her book, and on Tuesday,she read 2/6 of the book. How much has she read? A 1/6 B 1/5 C 1/3 D 1/2

OpenStudy (anonymous):

\[\frac{1}{6}+\frac{2}{6}=\frac{3}{6}\] is a start

OpenStudy (anonymous):

So, if it equals 3/6, divide each number by three. 3/3 = 1, 6/3 = 2, put them together as a fraction, and what do you get?

OpenStudy (xapproachesinfinity):

you only need to simplify @satellite73 give you the start

OpenStudy (ageta):

\[1/6 +2/6=3/6 =1/2\]

OpenStudy (ageta):

1/2 is the answer

OpenStudy (xapproachesinfinity):

hi no giving answers